Consider the following small open economy model with production. At dates 1 and 2, the home country receives exogenous fixed endowments y1 and y2 respectively. The home country has access to the international capital market at a fixed interest rate r* at which it can save or borrow. Let the net saving of the home country be s1 and its consumption stream in two periods be given by

c1 and c2 respectively.the following maximization problem:

Max ln c1 + ln c2


s.t. C1 + S1 = Y1


C2 = C1(1+r*) + Y2


Derive optimal consumption and current account functions and carefully interpret it in terms of a two-period Fisherian graph.


Refer to (a). Suppose the home country also has an access to an investment

technology which means that if it invests k units at date 1, it produces y2 = Aka units of output

in the next period where A >0 and 0 < a< 1 . Modify budget constraints for (a)Derive the optimal investment and saving rules for this

economy assuming the same logarithmic utility function as in (a). Interpret your results


1
Expert's answer
2021-05-20T13:51:47-0400

(i)

consumer will maximize utility when budget constraint is

"C_1+\frac{C_2}{1+r}=Y_1+\frac{Y_2}{1+r}"

solve the budget constraint for "C_2"

"C_2=Y_1(1+r)+Y_2-(1+r)C_1"

substitute the constraint into objective function

"U=u(C_1)+\u03b2(Y_1(1+r)+Y_2-(1+r)C_1)"

maximize the objective function wrt "C_1"

"U(C_1)-\u03b2u(C_2)(1+r)=0"


"U(C_1)\frac{1}{1+r}=β(C_2)"


"Bu(C_2), u(C_1)=\frac{1}{1+r}"

the expression shows how much "C_1" is willing to trade for one more unit with "C_2" given constant utility and also how much "C_1" have to give up to get one more unit "C_2" .


current account for any given r

"CA(r;Y_1;Y_2)=Y_1-C_1(r;Y*_1,Y*_2)"

Current account

"Y_1-C=\frac{Y_1-Y_2}{1+r}"


(ii)

optimal investment and saving rules

new budget constraint

"C_1+\frac{C_2}{1+r}=Y_1+\frac{Y_2}{1+r}+\frac{K_1-r(K_2-K_1)}{1+r}"

rule

"\frac{\delta m}{\delta r}=-\frac{Y_2-rK_2}{1-r_e}-\frac{K_2}{1+r}=-\frac{Y_2+K_2}{(1+r)_2}"

the slope is negative meaning investment and saving is decreasing with r as consumption increases.